# Towards Consistent and Controllable Image Synthesis for Face Editing

### v1

@@ Abstract
Current face editing methods mainly rely on GAN-based techniques, but recent focus has shifted to diffusion-based models due to their success in image reconstruction. However, diffusion models still face challenges in manipulating fine-grained attributes and preserving consistency of attributes that should remain unchanged. To address these issues and facilitate more convenient editing of face images, we propose a novel approach that leverages the power of Stable-Diffusion models and crude 3D face models to control the lighting, facial expression and head pose of a portrait photo. We observe that this task essentially involve combinations of target background, identity and different face attributes. We aim to sufficiently disentangle the control of these factors to enable high-quality of face editing. Specifically, our method, coined as RigFace, contains: 1) A Spatial Arrtibute Encoder that provides presise and decoupled conditions of background, pose, expression and lighting; 2) An Identity Encoder that transfers identity features to the denoising UNet of a pre-trained Stable-Diffusion model; 3) An Attribute Rigger that injects those conditions into the denoising UNet. Our model achieves comparable or even superior performance in both identity preservation and photorealism compared to existing face editing models.
@@ I Introduction
Lifelikely varying the lighting, expression, head pose, etc, of a portrait while keeping the identity and high-frequency facial characteristics has been a topic of enduring interest in face editing. This task requires complete disentangling and fine-grained control over identity, background, head pose and other face attributes. Current techniques for this mainly rely on GAN-based models [1, 2, 3]. However, the efficacy of GAN-based methods in editing real images is constrained by their dependence on GAN inversion to translate real images into a semantic latent space [4, 5, 6]. This can cause unintended change in the resulting images especially when dealing with face images that requires greater variations.
Recent research has sought to overcome these challenges by utilizing the potent generative potential of diffusion models. Diffusion Autoencoders (DAE) [7] has shown that diffusion models, when trained on autoencoding tasks, can create a latent space that facilitates appearance editing. It has the ability to interpolate between expressions, such as from smiling to a poker face, once semantic labels are applied to determine the direction of editing. Nevertheless, it can only edit with binary semantic labels, while 3D lighting, pose, and flexible expressions are hard to be expressed by only two directions. [8] instead leverages the power of large pre-trained diffusion models by introducing an adapter plugin. Although this can reduce the training cost, the generated outcomes are not satisfactory enough. Using adapters involves only a subset of parameters released, which can lead to the edited images closely resembling cartoons and struggling to deviate from the generative style of the original diffusion model.
@@
Meanwhile, Large diffusion models offer powerful feature representations by allowing us to leverage their knowledge and adapt them to specific tasks. This is achieved by inheriting their parameters and architectures while keeping all parameters trainable. Inspired by this, we are committed to developing an effective model that allows us to photorealisticly edit the head pose, lighting and facial expression of a given photo with the help of pre-trained latent diffusion models. The design motivation of our model is twofold: (1) Fully disentangling of the lighting, background, pose and facial expressions to enable consistent edit from three conditions; (2) Leveraging a large Stable-Diffusion model by inheriting its parameters as well as architecture to enable high-quality editing results.
Specifically, the proposed model comprises three components: 1) Spatial Attribute Provider is designed to automatically generate 3D renderings, the mask of dilated foreground area and expression parameters, which provides disentangled and precise guidance for controlled generation of lighting, head pose and facial expressions. Additionally, This strategy mitigates potential issues that may arise when extracting only the background from the identity image, such as inconsistencies caused by variations in the target background present in the training data. 2) Identity Encoder inherits a Stable-Diffusion model to extract face features and then transfers them to the transformer blocks of the denoising UNet. The way these features transmitted to the denoising UNet significantly enhances the consistency of identity in the generated images. 3) Attribute Rigger integrates all the conditions from identity to maintain background consistency, collect clues about global lighting and spatial reference, and generate facial expression in harmony with the expression parameters. Our model adheres to the given conditions, and meanwhile preserves most attributes of the identity image that are not intended to change. Contributions are summarized as follows:
Figure 2: Overview pipeline of our proposed method consisting three modules: 1) The Spatial Attribute Provider adapts the foreground mask and predict 3D rendering as well as expression coefficients, offering decoupled and more clear guidance for controlled generation. 2) The Identity Encoder improves identity consistency by transfering spatial features to the denoising UNet. 3) The Attribute Rigger integrates the image conditions into the denoising UNet to fill in the attributes aimed to edit.
  • We introduce RigFace, a facial editing model designed to facilitate precise control over head pose, lighting and facial expressions for a given identity. This model efficiently and proficiently enables face editing, surpassing previous state-of-the-art GAN-based and diffusion-based methods.
@@
  • RigFace is training-efficient by inheriting the knowledge from pre-trained latent diffusion models. By releasing all the parameters in the denoising UNet, it avoids having the style of generated images confined to those produced by stable-diffusion.
@@ II-B Facial Appearance Editing
Conditional generation and editing is a prominent area of research that focuses on guiding generative models using various modalities, such as text, segmentation masks, and audio. For instance, methods like StyleCLIP [13], DiffusionCLIP [14], and others have achieved impressive results in text-driven face generation and editing. Many derivative works in face editing, such as face swapping and reenactment [8, 15, 1, 16], often use 3D Morphable Face Models (3DMM) as a conditioning input for editing. 3DMMs offer a crucial parameter space for defining and resolving 3D facial features [17]. Some commonly used 3DMM derived from 4D scans like FLAME [18] and BFM [19] facilitate modifications in shape, pose, and facial expressions. While these models offers a concise and physically relevant space for facial geometry, it lacks capabilities for depicting appearance. Addressing this gap, some image reconstruction methods utilize 3DMM while also modeling facial appearance using Lambertian reflectance and Spherical Harmonics (SH) lighting. Trained on an extensive dataset, these models can deduce albedo, SH lighting, and FLAME parameters from just a single portrait photo. Inspired by these works, in this paper, we utilize two methods, i.e., DECA [20] and Deep3DFace [21] to provide expression, pose and lighting condtions that support the synthesis of new face by editing 3DMM parameters.
@@ III-C Spatial Attribute Provider
Specifically, we use a pre-trained face parsing model [37] to obtain the head region. It is important to note that directly applying the mask of the target image’s head to the identity image is impractical when no pose adjustment is required. To address this, we extract the head regions from both the source and target images, combine them into a new image, and then apply dilation to create a more adaptable area for the pose. The result image is then used to mask the foreground of the target image. To enable pose edit, we draw the face landmarks of the target using Deep3DRecon [21] on the masked area. This result in one of the conditions sent into Attribute Rigger in Fig. 2.
III-D Identity Encoder
In text-to-image tasks, the alignment between text and images is semantic, which indicates that most approaches using CLIP image encoder tend to extract high-level features of the identity. However, our task requires to keep detailed features of the identity image, so we design the identity encoder to enable consistent control.
We adopt the latent diffusion framework of Stable-Diffusion for the Identity Encoder, similar to the work of [38]. It only inherits the weights in the original model for initialization, and during the training all parameters are released. Given the source image $\mathbf{I}_{sor}$, a frozen VAE encoder $\mathcal{E}$ compresses the image into a low dimension feature representation and then sent it into the Indentity Encoder. The denoising UNet is also inherited from the latent diffusion model with all parameters released. As demonstrated in the rectangular box in the top left corner of Fig. 2, in each transformer blocks of these two models, suppose feature $\mathbf{x}_{id}\in\mathbb{R}^{h\times w\times c}$ from Identity Encoder and $\mathbf{x}_{dn}\in\mathbb{R}^{h\times w\times c}$ from denoising UNet, $\mathbf{x}_{id}$ is first concatenated with $\mathbf{x}_{dn}$ along the $w$ dimension. Then the self-attention layer takes in it and only half feature of the output along the $w$ dimension is kept and sent into the cross-attention layer. Note that the Identity Encoder only extracts features in the entire adding noise and denoising process, so in the inference stage it won’t cause extreme increase in time cost.
@@ III-E Attribute Rigger
Although ControlNet [11] has shown highly conditional generation capabilities for diffusion models, it needs a copy of the UNet which could induce a significant increase in computational complexity. Instead, we design the Attribute Rigger to be lightweight by utilizing a convolution layer with 8 channels, $4\times 4$ kernels and $2\times 2$ strides. To be specific, the VAE encoder $\mathcal{E}$ in Stable-Diffusion compresses the spatial conditions into latent features and then the Attribute Rigger takes in the features and integrates them into the denoising UNet $\boldsymbol{\phi}_{SD}$. The expression coefficients $\boldsymbol{\psi}$ are mapped by a linear layer to have the same dimension with the time embedding of the denoising UNet and then added to the time embedding. As a result, given all the conditions and identity features, the whole framework is optimized by minmizing the following:
@@ IV-A Experimental Setup
Dataset. During the training we leverage the Aff-Wild [39] dataset. We randomly sample 30K images as the source image, and accordingly sampled another 30K images in the same video as the target image. During the evaluation, we sampled 20 identities from the videos and operated 3 editings for expression, pose and lighting, which leads to a number of 180 generated images. The videos used in training and test are different to ensure reliability. For preprocessing of the sampled images, we follow FOMM [40] to crop the face, preserve partial backgrounds, and resize them into $512\times 512$ to fit in the resolution of Stable-Diffusion.
Evaluation Criteria. We use the ID feature similarity calculated by a pre-trained model  [^1] https://github.com/ageitgey/face_recognition to measure identity preservation. We use SSIM [41] to evaluate the background similarity between generated and real images. Note that for fair evaluation we mask the face and only preserve the background area for evaluation of this score. For expression edit evaluation, we compute the emotion discrepancy between features from the edited and source image as in EMOCA [42]. For pose edit evaluation, we compute the pose coefficients of FLAME [18] faces and then compute Root Mean Squared Error (RMSE). For lighting, we use DECA to infer the spherical harmonics and then compute RMSE.
@@ IV-B Experimental Results
Fig. 3 visually shows some example results of our method and other face editing methods on some identities. We observe that compared with other methods, our RigFace consistently achieves the best edit effects for various individuals and backgrounds. More specifically, 3DMM-based methods like DECA and HeadNerf often fails to preserve the background of the original image. The geometry, albedo, and lighting derived from 3DMM remain rudimentary and significantly distant from realistic representation. Although Deep3DRecon uses rendering to reattach the face to the original background image, when the head pose is edited, filling in the background becomes a challenge that rendering alone cannot resolve. GIF tends to change the identity and other facial attributes, despite its decent control over generating specific expressions, lighting, and poses. DiffusionRig alters certain facial features, and the overall image generation quality decreases. The face appears smoother compared with the input image. FaceAdapter performs better in preserving identity details, but some edits are entangled: it can cause expression variation when editing the lighting. In contrast, our model not only generates realistic images while preserveing most of the individual details, it also enable the decoupled editing of specific controls.
Tab. I presents the quantitative evaluation results on different methods. We find that all of the reported methods (except Deep3DRecon) do not show a significant performance gap in editing facial expressions and poses. DECA even performs the best in editing lighting and pose since it renders images directly using FLAME coefficients. Nonetheless, our method presents a significant improvement in preserving identity and background of the input image, as depicted by the large increase in ID and SSIM scores, which also corresponds well with the observations in the qualitative results.
@@
Generalization Ability. It is also worth highlighting that RigFace can generalize to out-of-domain identity images of unseen styles with unexpectedly strong control over appearance, even without additional fine-tuning on the target domain. Fig 4. compares the zero-shot results of applying HeadNerf [44], DiffusionRig [15], FaceAdapter [8] and RigFace to out-of-domain identities, whose visual style is distinct from corresponding training data of the real-human face images. We observe that most of existing methods are trained on real-human datasets, so we test our method on more comic characters. The comparison results show that RigFace generalizes surprisingly well to new character images even though it has never been trained on such data. Furthermore, we include more comic images as a small dataset which contains 1000 images by collecting them on a website  [^3] https://fi.pinterest.com/ to fine-tune our model, which we dub the small dataset as Comic-S. As shown in Tab. II, higher generation quality can also be achieved through fine-tuning on specific datasets.
@@
Comparison with Face Reenactment Methods. Since our method can be applied to edit pose, expression, and lighting together, it can effectively perform face reenactment. Fig. 5 compares with SoTA methods on face reenactment, including TPSM [45], FADM [16], HyperReenact [1], DiffusionRig [15], and FaceAdapter [8]. Tab. III reports the results tested on the same 20 identities for face reenactment task. Note that due to changes in pose, background computation may introduce interference for this task. Therefore, it is necessary to combine SSIM and pose metrics to evaluate background preservation. We can observe that we achieve comparable or even optimal results in image quality. Owing to the Attribute Rigger, integrating the target background area into the condition helps prevent interference from background motion. During inference, the model only needs to copy a large portion of background from the source, which significantly simplifies background generation and improves background consistency. More importantly, due to the design of the identity encoder, our model also performs well in preserving identity details.

### v2

@@ Abstract
Face editing methods, essential for tasks like virtual avatars, digital human synthesis and identity preservation, have traditionally been built upon GAN-based techniques, while recent focus has shifted to diffusion-based models due to their success in image reconstruction. However, diffusion models still face challenges in controlling specific attributes and preserving the consistency of other unchanged attributes especially the identity characteristics. To address these issues and facilitate more convenient editing of face images, we propose a novel approach that leverages the power of Stable-Diffusion (SD) models and crude 3D face models to control the lighting, facial expression and head pose of a portrait photo. We observe that this task essentially involves the combinations of target background, identity and face attributes aimed to edit. We strive to sufficiently disentangle the control of these factors to enable consistency of face editing. Specifically, our method, coined as RigFace, contains: 1) A Spatial Attribute Encoder that provides presise and decoupled conditions of background, pose, expression and lighting; 2) A high-consistency FaceFusion method that transfers identity features from the Identity Encoder to the denoising UNet of a pre-trained SD model; 3) An Attribute Rigger that injects those conditions into the denoising UNet. Our model achieves comparable or even superior performance in both identity preservation and photorealism compared to existing face editing models. Code is publicly available at https://github.com/weimengting/RigFace.
@@ I Introduction
Lifelikely varying the lighting, expression, head pose and other attributes of a portrait while keeping the identity and high-frequency facial characteristics has been a topic of enduring interest in face editing. Such capabilities are essential for applications in virtual avatars, digital human synthesis, and identity-preserving facial modifications. This task requires complete disentangling and fine-grained control over identity, background, head pose and other face attributes. Current techniques for this mainly rely on GAN-based models [1, 2, 3]. However, the efficacy of GAN-based methods in editing real images is constrained by their dependence on GAN inversion to translate real images into a semantic latent space [4, 5, 6]. This can cause unintended change in the resulting images especially when dealing with face images that requires greater variations.
Recent research has sought to overcome these challenges by utilizing the potent generative potential of diffusion models. However, these methods fail to efficiently preserve identity details and have not been successful in editing certain specific attributes. For instance, Diffusion Autoencoders (DAE) [7] has the ability to interpolate between expressions, such as from smiling to a poker face, once semantic labels are applied to determine the direction of editing. Nevertheless, it can only edit with binary semantic labels, while 3D lighting, head pose, and more flexible expressions are hard to be expressed by only two directions. Face-Adapter [8] instead leverages the power of large pre-trained diffusion models by introducing an adapter plugin. Although this can reduce the training cost, the generated outcomes are not satisfactory enough. Using adapters involves only a subset of parameters released, which can lead to the edited images resembling cartoons and struggling to deviate from the generative style of the original diffusion model, which are insufficiently close to lifelike images found in the real world. Some methods also proposed face editing plugins for large pre-trained diffusion models [9, 10]. However, these approaches mainly emphasize attribute editing through text, which inevitably compromises spatial control to maintain text-driven editability.
@@
To resolve the above challenges, we are committed to developing an effective model that allows us to photorealisticly edit the head pose, lighting and facial expression of a given photo with the help of pre-trained latent diffusion models. Our method is inspired the success of works that inherit the parameters and architectures while keeping all parameters trainable of Stable-Diffusion (SD) models, which has achieved better performance in adapting knowledge from SD models to specific tasks. The design motivation of our model is twofold: (1) Fully disentangling of the lighting, background, pose and facial expressions to enable consistent and independent edit from three conditions; (2) Leveraging a large SD model by inheriting its parameters as well as architecture to enable high-quality editing results.
Specifically, the proposed model, coined as RigFace, comprises three components: 1) Spatial Attribute Provider is designed to automatically generate 3D renderings, the mask of dilated foreground area and expression parameters, which provides disentangled and precise guidance for controlled generation of lighting, head pose and facial expressions. Additionally, this strategy mitigates potential issues that may arise when extracting only the background from the identity image, such as inconsistencies caused by variations in the target background present in the training data. 2) FaceFusion that transfers the face features from the Identity Encoder through layer-by-layer transmission of transformer blocks to the denoising UNet. The way these features transmitted to the denoising UNet significantly enhances the consistency of the identity details in the generated images. 3) Attribute Rigger integrates all the conditions from identity to maintain background consistency, collect clues about global lighting and spatial reference, and generate facial expression in harmony with the expression parameters. Our model adheres to the given conditions, and meanwhile preserves most attributes of the identity image that are not intended to change. Fig. 1 showcases some examples of the results edited by our model.
Our contributions are summarized as follows:
  • We introduce RigFace, a face editing model to facilitate precise control over head pose, lighting and facial expressions for a given identity. This model proficiently enables face editing, surpassing previous state-of-the-art GAN-based and diffusion-based methods.
@@
  • We propose FaceFusion to efficiently align the identity features with the target face in the self-attention of transformer blocks without redundant warping.
  • RigFace is powerful by inheriting the knowledge from pre-trained latent diffusion models. Releasing all parameters can unleash the full potential of the model, enabling the SD model to achieve better adaptability to specific tasks.
@@ II-B Facial Appearance Editing
Conditional generation and editing is a key research area that focuses on guiding generative models through various modalities, including text, segmentation masks, and audio. Notable methods such as StyleCLIP [15] and DiffusionCLIP [16] have demonstrated impressive results in text-driven face generation and editing. Beyond text-based approaches, face editing tasks like face swapping and reenactment [8, 17, 1, 18] often leverage 3D Morphable Face Models (3DMM) as conditioning inputs due to their well-defined and disentangled parameter space for representing 3D facial geometry [19]. Common 3DMMs, such as FLAME [20] and BFM [21], enable precise control over shape, pose, and expressions. However, while these models provide a compact and physically meaningful representation of facial structure, they lack the ability to model appearance of face images. To bridge this gap, some image reconstruction methods integrate 3DMMs with Lambertian reflectance and Spherical Harmonics (SH) lighting to capture facial appearance. Trained on large-scale datasets, these models can infer albedo, SH lighting, and FLAME parameters from a single portrait image, enhancing realism and editability [22, 23]. Inspired by these works, we employ 3D reconstruction to provide our model with precise conditioning on pose, expression and lighting, enabling accurate and controllable edits.
@@ III-A Preliminaries
Figure 2: Overview pipeline of our proposed method consisting three modules: 1) The Spatial Attribute Provider adapts the foreground mask and predict 3D rendering as well as expression coefficients, offering decoupled and more clear guidance for controlled generation. 2) The FaceFusion transfers spatial features of face from Identity Encoder to the denoising UNet to enable high identity consistency. 3) The Attribute Rigger integrates the image conditions into the denoising UNet to fill in the attributes aimed to edit.
@@ III-C Spatial Attribute Provider
Specifically, we use a pre-trained face parsing model [37] to obtain the head region. It is important to note that directly applying the mask of the target image’s head to the identity image is impractical when no pose adjustment is required. To address this, we extract the head regions from both the source and target images, combine them into a new image, and then apply dilation to create a more adaptable area for the pose. The result image is then used to mask the foreground of the target image. This result in one of the conditions sent into Attribute Rigger in Fig. 2.
III-D FaceFusion
To preserve the facial details of the identity image, we need to integrate the face information from the source image into the diffusion U-Net. Although many SD-based methods use CLIP as the reference image encoder, we find that it fails to preserve many detailed identity features, as validated in Sec. IV-C. Previous works [38, 39] have demonstrated that self-attention effectively preserves reference image content through information fusion. Inspired by this, we design FaceFusion to efficiently learn the detail features of the face characteristics from the Identity Encoder. The left side of Fig. 2 shows the architecture of the Identity Encoder, which is essentially identical to the denoising UNet of SD. It only inherits the weights in the SD model for initialization, and during the training all parameters are released. Given the source image $\mathbf{I}_{sor}$, a frozen VAE encoder $\mathcal{E}$ compresses the image into a low dimension feature representation and then sent it into the Indentity Encoder. The denoising UNet is also inherited from the SD model with all parameters released. As demonstrated in the FaceFusion blocks of Fig. 2, in each transformer blocks of these two models, suppose feature $\mathbf{x}_{id}\in\mathbb{R}^{h\times w\times c}$ from the Identity Encoder and $\mathbf{x}_{dn}\in\mathbb{R}^{h\times w\times c}$ from the denoising UNet, $\mathbf{x}_{id}$ is first concatenated with $\mathbf{x}_{dn}$ along the $w$ dimension. Then the self-attention layer takes in it and only half feature of the output along the $w$ dimension is kept and sent into the cross-attention layer. Note that the Identity Encoder only extracts features in the entire adding noise and denoising process, so in the inference stage it won’t cause extreme increase in time cost.
@@ III-E Attribute Rigger
Although ControlNet [13] has shown highly conditional generation capabilities for diffusion models, it needs a copy of the UNet which could induce a significant increase in computational complexity. Instead, we design the Attribute Rigger to be lightweight by utilizing a convolution layer with 8 channels, $4\times 4$ kernels and $2\times 2$ strides. To be specific, the VAE encoder $\mathcal{E}$ in SD compresses the spatial conditions into latent features and then the Attribute Rigger takes in the features and integrates them into the denoising UNet $\boldsymbol{\phi}_{SD}$. The expression coefficients $\boldsymbol{\psi}$ are mapped by a linear layer to have the same dimension with the time embedding of the denoising UNet and then added to the time embedding. As a result, given all the conditions and identity features, the whole framework is optimized by minmizing the following:
@@ IV-A Experimental Setup
Dataset. Since no dedicated dataset exists for training our task, we preprocess the Aff-Wild [40] video dataset to meet our training requirements. Aff-Wild is a large-scale in-the-wild dataset consisting of videos with a variety of subjects in different emotional states, head poses and illumination conditions. Each video segment typically features a single individual. We randomly sample 30K images from the videos as the source image, and accordingly sampled another 30K images in the same video as the target image. During the evaluation, we sampled 20 identities from the videos and operated 3 editings for expression, pose and lighting, which leads to a number of 180 generated images. The videos used in training and test are different to ensure reliability. For preprocessing of the sampled images, we follow FOMM [41] to crop the face, preserve partial backgrounds, and resize them into $512\times 512$ to fit in the resolution of Stable-Diffusion.
Evaluation Criteria. We use the ID feature similarity calculated by a pre-trained model  [^1] https://github.com/ageitgey/face_recognition to measure identity preservation. We use SSIM [42] to evaluate the background similarity between generated and real images. Note that for fair evaluation we mask the face and only preserve the background area for evaluation of this score. For expression edit evaluation, we compute the emotion discrepancy between features from the edited and source image as in EMOCA [43]. Following the work of [17], we use FLAME coefficients to evaluate the accuracy of edited pose and lighting. We compute the pose coefficients of FLAME [20] faces and then compute Root Mean Squared Error (RMSE) for pose edit evaluation. For lighting, we use DECA to infer the spherical harmonics and then compute RMSE.
@@ IV-B Experimental Results
Fig. 3 visually shows some example results of our method and other face editing methods on some identities. We observe that compared with other methods, RigFace consistently achieves the best edit effects for various individuals and backgrounds. Compared with 3DMM-based methods like DECA and HeadNerf, our method is able to preserve the background of the original image with minimal loss. Although Deep3DRecon uses rendering to reattach the face to the original background image, when the head pose is edited, filling in the background becomes a challenge that rendering alone cannot resolve. More importantly, our method accurately preserves and transfers facial details from the original image. GIF tends to change the identity and other facial attributes, despite its decent control over generating specific expressions, lighting, and poses. DiffusionRig alters certain facial features, and the overall image generation quality decreases. The face appears smoother compared with the input image. FaceAdapter performs better in preserving identity details, but some edits are entangled: it can cause expression variation when editing the lighting. In contrast, our model not only generates realistic images while preserveing most of the individual details, it also enable the decoupled editing of specific controls.
Tab. I presents the quantitative evaluation results on different methods. We can observe that our method presents a significant improvement in preserving identity and background of the input image, as depicted by the large increase in ID and SSIM scores. Although some of the 3DMM-based methods do not show a significant performance gap in editing facial expressions and poses since they render images directly using FLAME coefficients, our method surpasses them by achieving higher identity preservation, finer detail transfer, and more precise controllability.
@@
Generalization Ability. It is also worth highlighting that RigFace can generalize to out-of-domain identity images of unseen styles with unexpectedly strong control over appearance, even without additional fine-tuning on the target domain. Fig 4. compares the zero-shot results of applying HeadNerf [45], DiffusionRig [17], FaceAdapter [8] and RigFace to out-of-domain identities, whose visual style is distinct from corresponding training data of the real-human face images. We observe that most of existing methods are trained on real-human datasets, so we test our method on more comic characters. The comparison results show that RigFace generalizes surprisingly well to new character images even though it has never been trained on such data. Furthermore, we include more comic images as a small dataset which contains 1000 images by collecting them on a website  [^3] https://fi.pinterest.com/ to fine-tune our model, which we dub the small dataset as Comic-S. As shown in Tab. II, higher generation quality can also be achieved through fine-tuning on specific datasets, further enhancing identity consistency and editability based on the zero-shot capability.
@@
Comparison with Face Reenactment Methods. Since our method can be applied to edit pose, expression, and lighting together, it can effectively perform face reenactment. Fig. 5 compares with SoTA methods on face reenactment, including TPSM [46], FADM [18], HyperReenact [1], DiffusionRig [17], and FaceAdapter [8]. Tab. III reports the results tested on the same 20 identities for face reenactment task. Note that due to changes in pose, background computation may introduce interference for this task. Therefore, it is necessary to combine SSIM and pose metrics to evaluate background preservation. We can observe that we achieve comparable or even optimal results in identity consistency and editing precision. Owing to the Attribute Rigger, integrating the target background area into the condition helps prevent interference from background motion. During inference, the model only needs to copy a large portion of background from the source, which significantly simplifies background generation and improves background consistency.
Figure 6: Ablation study for different colors used to mask for the background image. Black foreground is easy to clash with the color of the original images.